Find the equation of a circle given its center C and one point P.
C(0,4),P(0,0).
Thank You

Answer by scott8148(6628)

general circle equation ___ (x - h)^2 + (y - k)^2 = r^2 ___ center is (h,k), radius is r

(x - 0)^2 + (y - 4)^2 = 4^2